OpenStudy (anonymous):

simplify the following expression. show your work [x/(x-4)+6/(x+4)]/3/(2x^2-32)

OpenStudy (campbell_st):

well simplify the numerator by having a common denominator \[\frac{x}{x -4} + \frac{6}{x+4} = \frac{x(x+4) + 6(x -4)}{(x+4)(x - 4}\] you need to simplify the numerator the denominator of your algebraic faction can be simplified to \[\frac{3}{2(x-4)(x+4)}\] find the reciprocal of the denominator and multiply \[\frac{x(x+4) +6(x-4)}{(x+4)(x-4)} \times \frac{2(x-4)(x+4)}{3}\] just cancel common factors and simplify

OpenStudy (anonymous):

thank you so much